OpenStudy (anonymous):

determine convergence of divergence: (2/5)+(2*6)/(5*8)+(2*6*10)/(5*8*10)+(2*6*10*14)/(5*8*11*14)+...

OpenStudy (psymon):

Just looking at the pattern, Im hoping there was a slight typo

OpenStudy (anonymous):

damn i wrote it wrong....the third term on the demoniator should be (2*6*10)/(5*8*11)

OpenStudy (psymon):

Bleh, just need a way to turn this into a series in terms of n.

OpenStudy (anonymous):

well since the first term is (2/5) can we factor that out some how? ANd it also looks like a factorial with something being added....Idk

OpenStudy (psymon):

Yeah, Ive been messing with factorials, too. As far as the 2/5, it just makes me think that it's left over due to n = 0 making things become 1. x^0 or (0)! could cause that. So trying to get ideas of how to mess with that.

OpenStudy (psymon):

Makin my head hurt, haha. It just looks like introduction to power series, but brain isnt coming up with anything.

OpenStudy (psymon):

I'll stick around to see if someone can point us in the right direction, but brain needs to relax xD

OpenStudy (anonymous):

totally undertand...thanks for your help anyways !!!

OpenStudy (anonymous):

Your series diverges. Each term is the previous term times a fraction. After the fourth term these multulpliers become \[ \frac {14} {14}, \ \frac {18}{17}, \frac {22}{20}, \frac {26}{23}, \frac {30}{26} \] The sequence of n-terms become increasing, so it cannot go to zero. So the series is divergent

OpenStudy (anonymous):

\[ a_1=\frac 2 5\\ a_2 = a_1\frac 6 8\\ a_3 = a_2\frac {10}{ 11}\\ a_4 = a_3\frac {14}{ 14}=a_3\\ a_5 = a_4\frac {18}{ 17}> a_4\\ a_6 = a_5\frac {22}{ 20}>a_5\\ \cdots \]

OpenStudy (anonymous):

@Psymon

OpenStudy (anonymous):

Elias thank you very much for your help!!!

OpenStudy (psymon):

Is there a test that you can use to prove the divergence, though? I can see that it will diverge by regular analysis, but I would think that you'd have to say you used some test, right?

OpenStudy (anonymous):

Yes, this is called the divergence test. If \( lim_{n->\infty} a_n \ne 0\) then \[ \sum_1^{\infty} a_n \] is divergent.

OpenStudy (psymon):

And what you were showing was proof enough of that? I mean, I see it proves it, but Id be paranoid about whether or not it was considered rigorous enough I suppose.

OpenStudy (anonymous):

Believe me, it is an excellent proof of it. I showed that \( lim_{n\to \infty} a_n \ne 0 \) and that is enough to show that the series is not convergent. See http://www.math.oregonstate.edu/home/programs/undergrad/CalculusQuestStudyGuides/SandS/SeriesTests/divergence.html